Part A: What is the concentration of K^+ in 0.15 M of K2S ?

Part B: If \rm CaCl_2 is dissolved in water, what can be said about the concentration of the \rm Ca^{2+} ion?
It has the same concentration as the \rm Cl^- ion.
Its concentration is half that of the \rm Cl^- ion.
Its concentration is twice that of the \rm Cl^- ion.
Its concentration is one-third that of the \rm Cl^- ion.
Part C: A scientist wants to make a solution of tribasic sodium phosphate, \rm Na_3PO_4, for a laboratory experiment. How many grams of \rm Na_3PO_4 will be needed to produce 700 mL of a solution that has a concentration of \rm Na^+ ions of 0.600 \it M?
Express your answer numerically in grams.
Answer #1
Concepts and reason

A balance chemical equation of a reaction is the simple representation of the reaction, which gives the idea of the particles and participates in the reaction. Concentration is the amount of substance present in a given volume of a solution.

Mole is the mass that explains the particles present in the substance.

Fundamentals

A balance chemical equation of a reaction provides the number of particles and participates in a chemical reaction.

The molarity of a solution is the number of moles of the solute present in 1L of the solution.

Mole is the mass that explains the particles present in the substance. A mole can be determined by using given molarity and volume.

Numberofmoles=GivenmassingramMolecularmass{\rm{Number}}\,{\rm{of}}\,{\rm{moles}}\,{\rm{ = }}\,\frac{{{\rm{Given}}\,{\rm{mass}}\,{\rm{in}}\,{\rm{gram}}}}{{{\rm{Molecular}}\,{\rm{mass}}}}\,

Numberofmoles=Molarity×VolumeinL{\rm{Number}}\,{\rm{of}}\,{\rm{moles}}\,{\rm{ = }}\,{\rm{Molarity \times }}\,{\rm{Volume}}\,{\rm{in}}\,{\rm{L}}\,

(A)

Balanced chemical equation of the reaction is

K2S(aq)2K+(aq)+S2(aq){{\rm{K}}_{\rm{2}}}{\rm{S}}\left( {{\rm{aq}}} \right)\,\longrightarrow{{}}\,{\rm{2}}{{\rm{K}}^{\rm{ + }}}\left( {{\rm{aq}}} \right)\,{\rm{ + }}\,{{\rm{S}}^{{\rm{2 - }}}}\left( {{\rm{aq}}} \right)

From the chemical equation, one mole of K2S{{\rm{K}}_{\rm{2}}}{\rm{S}} produces two moles of K+{{\rm{K}}^{\rm{ + }}}

Therefore,

ConcentrationofK+=2×.15=0.30M\begin{array}{c}\\{\rm{Concentration}}\,{\rm{of}}\,{{\rm{K}}^{\rm{ + }}}\,{\rm{ = }}\,{\rm{2 \times }}{\rm{.15}}\\\\{\rm{ = }}\,{\rm{0}}{\rm{.30M}}\\\end{array}

(B)

Balanced chemical equation,

CaCl2(aq)Ca2+(aq)+2Cl(aq){\rm{CaC}}{{\rm{l}}_{\rm{2}}}\left( {{\rm{aq}}} \right)\,\longrightarrow{{}}\,{\rm{C}}{{\rm{a}}^{{\rm{2 + }}}}\left( {{\rm{aq}}} \right)\,{\rm{ + 2C}}{{\rm{l}}^{\rm{ - }}}\,\left( {{\rm{aq}}} \right)

From the equation,

[Ca2+]=12[Cl]\left[ {{\rm{C}}{{\rm{a}}^{{\rm{2 + }}}}} \right]\,{\rm{ = }}\,\frac{{\rm{1}}}{{\rm{2}}}\left[ {{\rm{C}}{{\rm{l}}^{\rm{ - }}}} \right]

Therefore, concentration of the Ca2+{\rm{C}}{{\rm{a}}^{{\rm{2 + }}}} ion is half of the concentration of Cl{\rm{C}}{{\rm{l}}^{\rm{ - }}} ion.

(C)

Numberofmoles=Molarity×VolumeinL{\rm{Number}}\,{\rm{of}}\,{\rm{moles}}\,{\rm{ = }}\,{\rm{Molarity \times }}\,{\rm{Volume}}\,{\rm{in}}\,{\rm{L}}\,

molesofNa+ions=600×700×103=0.420\begin{array}{c}\\{\rm{moles}}\,{\rm{of}}\,{\rm{N}}{{\rm{a}}^{\rm{ + }}}\,{\rm{ions}}\,{\rm{ = }}\,{\rm{600 \times 700 \times 1}}{{\rm{0}}^{{\rm{ - 3}}}}\\\\\,{\rm{ = }}\,{\rm{0}}{\rm{.420}}\\\end{array}

Balanced chemical equation,

Na3PO4(aq)3Na+(aq)+PO43(aq){\rm{N}}{{\rm{a}}_{\rm{3}}}{\rm{P}}{{\rm{O}}_{\rm{4}}}\left( {{\rm{aq}}} \right)\,\longrightarrow{{}}\,{\rm{3N}}{{\rm{a}}^{\rm{ + }}}\left( {{\rm{aq}}} \right)\,{\rm{ + }}\,{\rm{PO}}_{\rm{4}}^{{\rm{3 - }}}\left( {{\rm{aq}}} \right)

From the equation, one mole of Na3PO4{\rm{N}}{{\rm{a}}_{\rm{3}}}{\rm{P}}{{\rm{O}}_{\rm{4}}} produces three moles of Na+{\rm{N}}{{\rm{a}}^{\rm{ + }}} ions.

Therefore,

molesofNa3PO4=0.4203=0.140molecularmassofNa3PO4=163.94gmassofNa3PO4=0.140×163.94=22.95g\begin{array}{c}\\{\rm{moles}}\,{\rm{of}}\,{\rm{N}}{{\rm{a}}_{\rm{3}}}{\rm{P}}{{\rm{O}}_{\rm{4}}}\,{\rm{ = }}\,\frac{{{\rm{0}}{\rm{.420}}}}{{\rm{3}}}\\\\{\rm{ = }}\,{\rm{0}}{\rm{.140}}\\\\{\rm{molecular mass of N}}{{\rm{a}}_{\rm{3}}}{\rm{P}}{{\rm{O}}_{\rm{4}}}\,{\rm{ = }}\,{\rm{163}}{\rm{.94g}}\\\\{\rm{mass}}\,{\rm{of}}\,{\rm{N}}{{\rm{a}}_{\rm{3}}}{\rm{P}}{{\rm{O}}_{\rm{4}}}\,{\rm{ = }}\,{\rm{0}}{\rm{.140 \times 163}}{\rm{.94}}\\\\{\rm{ = }}\,{\rm{22}}{\rm{.95}}\,{\rm{g}}\\\end{array}

Ans: Part A

ConcentrationofK+=0.30M{\rm{Concentration}}\,{\rm{of}}\,{{\rm{K}}^{\rm{ + }}}\,{\rm{ = }}\,{\rm{0}}{\rm{.30M}}

Part B

It has the same concentration as the cl ion
Its concentration is half that of the Cl ion
Its concentration is twice that of

Part C

massofNa3PO4=22.95g{\rm{mass}}\,{\rm{of}}\,{\rm{N}}{{\rm{a}}_{\rm{3}}}{\rm{P}}{{\rm{O}}_{\rm{4}}}{\rm{ = }}\,{\rm{22}}{\rm{.95}}\,{\rm{g}}